# Symbolic Control for Autonomous Docking of Marine Surface Vessels

### v1

@@ Abstract
Docking marine surface vessels remains a largely manual task due to its safety-critical nature. In this paper, we develop a hierarchical symbolic control architecture for autonomous docking maneuvers of a dynamic positioning vessel, to provide formal safety guarantees. At the upper-level, we treat the vessel’s desired surge, sway, and yaw velocities as control inputs and synthesize a symbolic controller in real-time. The desired velocities are then transmitted to and executed by the vessel’s low-level velocity feedback control loop. Given a synthesized symbolic controller, we investigate methods to optimize the performance of the proposed control scheme for the docking task. The efficacy of this methodology is evaluated on a low-fidelity simulation model of a marine surface vessel in the presence of static and dynamic obstacles and, for the first time, through physical experiments on a scaled model vessel.
Index Terms:
Marine Robotics, Hybrid Logical/Dynamical Planning and Verification, Robot Safety, Correct-by-construction Synthesis, Symbolic Controllers
@@ I Introduction
In recent decades, there has been an increasing demand for enhancing the efficiency and operational safety of marine surface vessels (MSV) through the development of autonomy [1]. While low-level automatic control functionalities, such as dynamic positioning [2] are well studied, significant challenges remain in high-level control tasks like automatic path planning, collision avoidance, and guidance, as they are typically complex logical tasks and the corresponding controller synthesis over continuous (hybrid) systems is challenging. Traditionally, these tasks have been performed by operators who rely heavily on training, experience, situational awareness, and risk understanding. In contrast, autonomy leverages advanced sensing, computation, and control capabilities, offering the potential to enhance safety, improve usage of human resources, reduce environmental impact, and minimize human injuries and fatalities at sea.
Docking is among the various complex high-level control tasks for an MSV. When a vessel approaches a harbor for docking, it first enters the harbor basin where it must safely maneuver to a docking location while avoiding structures and other vessels; this is often referred to as the approach phase. After approaching the target docking location, the vessel’s auto-docking system will ensure the vessel is brought slowly and safely to berth. Docking is a highly complex maneuver that largely remains manually performed due to the high risk of collision and strict precision requirements  [3]. In this paper, we focus on the approach phase of the docking operation. Specifically, we aim to design a correct-by-construction control scheme for safe autonomous docking using rigorous tools from control theory and formal methods [4]. These tools, originating in computer science, are intended to ensure the correct operation of computer programs and digital circuits. Further, they offer formal guarantees, allowing us to derive controllers in a provable fashion. The correct-by-construction methodology, an emerging technique for safe autonomy over the past decade, represents a significant improvement over the common practice of verification after design completion [5], which typically necessitates extensive testing and re-design cycles.
@@
Figure 1: Symbolic control scheme for Dynamic Positioning (DP) vessels with low-level velocity feedback control loop.
@@
In the context of correct-by-construction synthesis, abstraction-based approaches (a.k.a. symbolic control) have been explored for various systems, including linear systems [6], nonlinear systems with bounded disturbances [7], and nonlinear stochastic games [8]. Specifically, a finite abstraction (symbolic system) with finite states and input sets is first constructed as a simplified version of the original continuous system. Controllers are then synthesized over the symbolic system to meet complex logical high-level control objectives and subsequently refined back to the original system to achieve formal safety guarantees. However, because the finite state and input sets are created by partitioning the original sets with grids, applying these approaches to real-world applications is challenging due to their computational complexity.
Contribution: In this paper, we propose a hierarchical symbolic control scheme to address the safe autonomous docking problem for Dynamic Positioning (DP) vessels, potentially extending to more complex logical control objectives, as illustrated in Fig. 1. At the upper-level, leveraging the vessel’s kinematic model with desired surge, sway, and yaw velocities as control inputs, a symbolic controller is synthesized in real-time using the approach in [9]. The desired velocities are then transmitted to and executed by a low-speed velocity feedback control loop in the DP control system, at the lower-level. To achieve real-time synthesis of the symbolic controller, we deploy pFaces [10] on a server to enable parallel computation for the synthesis procedure with graphics processing units (GPUs), thus mitigating the computational bottleneck of symbolic control methodologies. Furthermore, given a synthesized symbolic controller, we investigate methods to optimize performance for safe autonomous harbor maneuvering. The effectiveness of the proposed control scheme is validated on a low-fidelity simulation model of a MSV and, for the first time, through physical experiments on a scaled model vessel.
Related Works: Although docking is mostly performed manually due to its safety-critical nature, autonomous docking has gained increasing attention over the past decades [3]. Existing autonomous docking methods typically deploy various path planning algorithms, such as Dubins curves [11], A[12], and State-lattice[13], combined with accurate navigation systems and electronic nautical charts. The paths are then tracked using different control methodologies, including PID [14], adaptive control [15], model predictive control [16], and artificial neural networks [17]. However, due to the heuristic nature of many of these approaches and the safety-critical aspect of docking, the corresponding control software must be verified to ensure safety levels at least equivalent to those of human-navigated vessels [18]. Additionally, these approaches still have limitations in handling dynamic obstacles and situational awareness [3]. In contrast, the symbolic control scheme we use here provides inherent safety guarantees. These formal guarantees are achieved through the correct-by-construction synthesis paradigm, and the proposed real-time controller synthesis mechanism allows adaptation to dynamic obstacles and environments. While [19] also utilized correct-by-construction synthesis for safe autonomous harbor maneuvering, the controller was synthesized offline, and no physical experiments were conducted to validate the proposed results. Although symbolic control was hampered by computational complexity in the past, real-time synthesis is now becoming possible with parallel computation and GPUs  [20, 21]. This paper contributes to the ongoing efforts and demonstrates symbolic control on autonomous docking maneuvers with experiments on a scale marine vessel.
Organization: The system dynamics are introduced in Section II. In Section III, a hierarchical symbolic control strategy is utilized to synthesize and implement a control law for safe autonomous docking of a vessel. The experimental setup is introduced in Section IV, followed by the simulation and experimental results in Section V.
@@ II System Dynamics
The motion of surface vessels has been extensively studied; see e.g  [22]. In this project, we use a three degrees of freedom (3-DOF) maneuvering ship model that considers only surge, sway, and yaw motion. The vessel’s dynamics are modeled in a body-fixed frame with respect to a world-fixed frame, as shown in Fig. 2.
$\eta_{\psi}$(yaw)$x_{b},\nu_{x}$ (surge)$y_{b},\nu_{y}$ (sway)NorthEast$(\eta_{x},\eta_{y})$$x_{n}$$y_{n}$
Figure 2: Coordinate frame of marine vessel control system. The vessel is equipped with two azimuth thrusters and one tunnel thruster. The vessel’s position is described by vector $\eta$. The subscript $n$ refers to the world-fixed frame, while the subscript $b$ refers to the body-fixed frame.
II-A Kinematics
The kinematic model describes the relationship between the vessel’s body-fixed velocities and its position and orientation in the world-fixed frame. This relationship is expressed as
@@ II-A Kinematics
where $\eta=[\eta_{x},\eta_{y},\eta_{\psi}]^{\top}$ represents the vessel’s position $(\eta_{x},\eta_{y})$ and orientation $(\eta_{\psi})$ in the world-fixed frame. Furthermore, $\nu=[\nu_{x},\nu_{y},\nu_{\psi}]^{\top}$ denotes the vessel’s linear and angular velocities in the body-fixed frame, corresponding to surge, sway, and yaw motions, respectively. $\mathbf{R}(\eta_{\psi})$ is the rotation matrix that transforms velocities from the body-fixed frame to the world-fixed frame. In this work, linear velocities are expressed in meters per second ($m/s$) and angular velocities in radians per second ($rad/s$), unless stated otherwise.
II-B Kinetics
@@
where $\mathbf{M}\in\mathbb{R}^{3\times 3}$ is the inertia matrix, including the vessel’s mass and added mass terms, $\mathbf{C}(\nu)\in\mathbb{R}^{3\times 3}$ represents Coriolis and centripetal forces, $\mathbf{D}(\nu)\in\mathbb{R}^{3\times 3}$ is the hydrodynamic damping matrix accounting for drag forces, $\tau\in\mathbb{R}^{3}$ represents the control forces and moments applied to the vessel (e.g., thrusters), and $b\in\mathbb{R}^{3}$ represents unmodeled dynamics and slowly varying environmental loads. From (2), the vessel’s acceleration in terms of the control input $\tau$ is:
$\dot{\nu}=\mathbf{M}^{-1}(\tau-\mathbf{C}(\nu)\nu-\mathbf{D}(\nu)\nu+b).$ (3)
@@ III Control Strategy
The goal is to design a correct-by-construction hierarchical control architecture for a surface vessel that synthesizes a symbolic controller in real-time to enhance performance in autonomous harbor and docking maneuvers while avoiding obstacles. In Fig. 1, we highlight the upper-level and lower-level controllers of this architecture. We provide an explanation of symbolic control in Section III-A to describe the formation of the upper-level controller, and in Section III-B we elaborate on the low-level velocity control loop.
@@ III-A Symbolic Control
Symbolic control [7, 23, 24, 25] is an approach to derive provably safe controllers that satisfy complex logical specifications including objectives and constraints in linear temporal logic, a tool for reasoning about propositions qualified in terms of time. Symbolic control relies on the existence of a symbolic system which is a simplified, finite abstraction of the continuous system. We utilize the approach presented in [7] to obtain an abstraction of our system through a feedback refinement relation that ensures our symbolic model contains all of the behaviors of the vessel’s physical model.
The closed loop resulting from this approach can be seen in Fig. 3. We create a symbolic controller that consists of a quantizer and an abstract controller synthesized over the finite abstraction that is usually computed in the form of a finite automaton whose states are a result of the quantized plant model. The symbolic controller outputs a list of possible control actions that satisfy a given specification. We optimize over this list to determine the optimal control input for our system, a vessel. This results in a velocity command that is given back to the vessel and executed by a low-level velocity controller, as shown in Fig. 1.
@@
Figure 3: Closed loop resulting from the abstraction, refinement, and optimization approach, adapted from [7, Fig. 1].
Let us consider the control problem associated with the autonomous harbor and docking maneuvers from our two obstacle real-world experiments. We adopt the convention that the states of system $i$ are denoted with subscript $i$ and $[a,b]$ denotes a discrete interval. We seek to guide the vessel (1) throughout a region with boundary, $B$, into a target docking position, $T_{1}$, while avoiding obstacles $O^{1}_{1}$ and $O^{2}_{1}$ as seen in Fig. 4. We measure the state of the vessel and provide control inputs every $2$ seconds. This formulates our control problem $(S_{1},\Sigma_{1})$ where $S_{1}$ is a mathematical representation of the physical vessel model and $\Sigma_{1}$ is the following specification
@@
with $U_{1}$ = $[-0.2,0.2]\times[-0.2,0.2]\times[-0.3,0.3]$ as the set of inputs, and $X_{1}=\mathbb{R}^{2}\times[-\pi,\pi)$ as the set of vessel states.
To synthesize an abstract controller, we need to form a finite-state approximation of $S_{1}$. The resulting abstraction, $S_{2}$, is built using a feedback refinement relation. Intuitively, every state of $S_{2}$ will be associated to a subset of the state set of $S_{1}$ based on the refinement relation, and $S_{2}$ will have imposing relations among its states to represent the dynamics of the original system. These relations are defined as a transition function, as discussed below. In a general sense, $S_{2}$ will fully contain the behavior of $S_{1}$; however, $S_{2}$ is a finite-state approximation of our control problem over which the control synthesis problem can be solved using standard search algorithms.
[…]
@@
$S_{2}$ is connected to the plant via a static quantizer, irrespective of the specification imposed on the plant. Additionally, in defining $S_{2}$, we introduce margins for over-approximations of obstacles and an under-approximation for the target docking position. This can be seen in Fig. 4 as $O^{1}_{2}$, $O^{2}_{2}$, and $T_{2}$. These margins account for the geometrical shape of the vessel. Alternatively, they can be set according to other factors, e.g.  [26, 27]. Therefore, the abstract specification, $\Sigma_{2}$, of our system is (4) where we substitute $O^{1}_{1},O^{2}_{1},T_{1}$ with $O^{1}_{2},O^{2}_{2},T_{2}$, respectively.
@@
The abstract system, $S_{2}$, has an associated transition function that defines the relationship amongst its states. This transition function represents the plant behavior and determines where in the state space the plant may transition. Fig. 5 illustrates a transition function. A standard search algorithm [24] is performed to determine safe control commands that allow the vessel to achieve it’s goal position. A list of all possible control actions that solve this problem are returned each time the synthesis task is performed with an updated vessel position.
Once a list of safe possible actions is computed, the plant may be given any one of the control commands. Every action on this list will result in safe progress towards the plant’s goal. We present an approach to optimize this selection process for realistic vessel behavior. The optimal control action is chosen using a cost function that favors forward motion and reduces the norm between sequential control actions. Let $\mathcal{A}=\{\sigma_{1},\sigma_{2},\dots,\sigma_{n}\}$ be a list of actions where $\sigma_{i}\in\mathbb{R}^{3}$, $\sigma_{pre}$ be the previously chosen action, and $\Delta\sigma_{i}:=\sigma_{i}-\sigma_{pre}$. We define the cost function $J(\sigma_{i})$ as
@@
where $W\in\mathbb{R}^{6\times 6}$ is positive semi-definite. We select the control action that minimizes the cost
$\nu_{ref}:=\sigma^{*}=\text{arg}\min_{\sigma_{i}\in\mathcal{A}}J(\sigma_{i}).$ (6)
@@ III-B Velocity Feedback Controller
The velocity feedback controller, illustrated in Fig. 6, controls the vessel’s surge, sway, and yaw velocities, represented by $\nu_{ref}$, using a Multiple-Input Multiple-Output (MIMO) PID velocity controller, an extended Kalman filter (EKF) for state estimation, and a maneuvering based dynamic thrust allocation algorithm [28]. The PID controller calculates the desired force, $\tau_{cmd}$, to achieve the reference velocity $\nu_{ref}$, and the thrust allocation algorithm then determines the necessary thruster forces (see Fig. 8), $F_{1},F_{2},F_{3}$, and azimuth angles, $\alpha_{1},\alpha_{2}$, to apply these commands. Position and velocity predictions are computed using the robot_localization package’s EKF implementation, as described in [29]. The measurement sources vary based on the setup, as detailed in Section IV.

### v2

@@ Abstract
We develop a hierarchical control architecture for autonomous docking maneuvers of a dynamic positioning vessel and provide formal safety guarantees. At the upper-level, we treat the vessel’s desired surge, sway, and yaw velocities as control inputs and synthesize a symbolic controller in real-time. The desired velocities are then executed by the vessel’s low-level velocity feedback control loop. We next investigate methods to optimize the performance of the proposed control scheme. The results are evaluated on a simulation model of a marine surface vessel in the presence of static obstacles and, for the first time, through physical experiments on a scale model vessel.
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Autonomy is critical for enhancing the efficiency and operational safety of marine surface vessels (MSV) [1]. Low-level automatic control functionalities, such as dynamic positioning [2], are well studied, but significant problems remain in providing formal guarantees for high-level control tasks like automatic path planning, collision avoidance, and guidance. Traditionally, these tasks have been performed by operators who rely on training, experience, situational awareness, and risk understanding. However, in 2023, the European Maritime Safety Agency, reported a total of 418 vessel collision occurrences, where human actions were identified as the primary cause, accounting for 58.4% of the total [3]. In contrast, autonomy leverages advanced sensing, computation, and control capabilities, offering enhanced safety, improved usage of human resources, reduced environmental impact, and minimized human injuries and fatalities.
@@
Figure 1: Symbolic control architecture for Dynamic Positioning (DP) vessels.
Docking is among the complex high-level control tasks for MSVs, which largely remains manually performed due to the high risk of collision and strict precision requirements [4]. When a vessel approaches a harbor for docking, it first enters the harbor basin where it must safely maneuver to a docking location while avoiding structures and other vessels. Existing autonomous docking methods use various planning algorithms, such as Dubins curves [5], A[6], state-lattice[7], and Voronoi diagrams [8], combined with navigation systems, electronic nautical charts, and low-level control methods, including PID [9], adaptive control [10], linear quadratic regulator [11], model predictive control [12], and artificial neural networks [13]. However, due to the heuristic nature of many of these approaches, the safety of docking cannot be verified a priori.
To address the unique safety challenges, we propose a correct-by-construction control scheme using rigorous tools from control theory and formal methods [14]. These tools, originating in computer science, are intended to ensure the correct operation of computer programs and digital circuits. Further, they offer formal, or mathematical, guarantees of correct behavior with respect to a formal system specification a priori—a significant improvement over the common practice of verification after design completion [15], which typically necessitates extensive testing and redesign cycles [16].
@@
Among correct-by-construction synthesis methods, symbolic control “abstracts” the original continuous system into a symbolic system with finite states and input sets. Controllers are then synthesized over the symbolic system to meet complex logical high-level control objectives and subsequently refined back to the original system. Since the finite abstraction covers all behaviors of the underlying dynamical system, this approach gives formal safety guarantees by construction. Symbolic control has been explored for linear systems [17], nonlinear systems with bounded disturbances [18], and nonlinear stochastic games [19].
In this paper, we propose a hierarchical, real-time symbolic control scheme to address the safe autonomous docking problem for Dynamic Positioning (DP) vessels and present real-world experiments. At the upper-level, we leverage the vessel’s kinematic model with desired surge, sway, and yaw velocities as control inputs, and synthesize a symbolic controller in real-time using the approach in [20]. At the lower-level, the desired velocities are executed by a low-speed velocity feedback control loop in the DP control system, assigning control signals to the actuators of the vessel. As the vessel moves to the docking pose, it continuously updates obstacle positions and re-sythesizes the symbolic controller, enabling the symbolic controller to make informed decisions.
A traditional shortcoming of symbolic control is that the finite state and input sets are created by partitioning the original sets with grids, adding to the computational complexity of this method and hampering practical applications. To overcome this limitation and to enable real-time synthesis, we employ parallel computation with graphics processing units (GPUs) on a server running pFaces [21]. We then identify the optimal control input from the synthesized action list. The effectiveness of the proposed scheme is demonstrated on a simulation model of a MSV and — for the first time — through physical experiments on a scale model vessel. Additionally, we show that we can synthesize and re-synthesize a symbolic controller in real time, which is critical for situational awareness in a dynamic environment. Experiments in a scaled environment are an important step in tuning and deploying software on real-world vessels [22]; it allows us to close the sim-to-real gap efficiently and safely.
In earlier work [23], we investigated correct-by-construction synthesis for safe autonomous harbor maneuvering, but the method was not amenable to real-time synthesis and no physical experiments were conducted. The contributions of the present paper are: 1) to leverage parallel computation and GPUs as proposed in [24, 25] to make real-time synthesis possible, allowing us to re-synthesize a symbolic controller in the presence of a dynamic environment, and 2) to demonstrate the results experimentally — a crucial step towards deployment in real-world autonomous docking. The paper marks a new stage of maturity for symbolic control, demonstrating real-time synthesis capability, on a practical problem of great interest to the maritime industry.
@@ II System Dynamics
We use a three degrees of freedom (3-DOF) maneuvering ship model that considers surge, sway, and yaw motion [26]. The kinematic model describes the relationship between the vessel’s body-fixed velocities and its position and orientation in the world-fixed frame. This relationship is expressed as
@@
where $\eta=[\eta_{x},\eta_{y},\eta_{\psi}]^{\top}$ represents the vessel’s position $(\eta_{x},\eta_{y})$ and orientation $(\eta_{\psi})$ in the world-fixed frame. Furthermore, $\nu=[\nu_{x},\nu_{y},\nu_{\psi}]^{\top}$ denotes the vessel’s linear and angular velocities in the body-fixed frame, corresponding to surge, sway, and yaw motions, respectively. $\mathbf{R}(\eta_{\psi})$ is the rotation matrix that transforms velocities from the body-fixed frame to the world-fixed frame. In this work, linear velocities are expressed in meters per second ($m/s$) and angular velocities in radians per second ($rad/s$).
@@
where $\mathbf{M}\in\mathbb{R}^{3\times 3}$ is the inertia matrix, including the vessel’s mass and added mass terms, $\mathbf{C}(\nu)\in\mathbb{R}^{3\times 3}$ represents Coriolis and centripetal forces, $\mathbf{D}(\nu)\in\mathbb{R}^{3\times 3}$ is the hydrodynamic damping matrix accounting for drag forces, $\tau\in\mathbb{R}^{3}$ represents the control forces and moments, and $b\in\mathbb{R}^{3}$ represents slowly varying environmental or vessel loads.
@@ III Control Strategy
We propose a hierarchical control architecture for a surface vessel that synthesizes a symbolic controller in real-time to enhance performance in autonomous docking maneuvers. Fig. 1 highlights the upper- and lower-level controllers of this architecture. Section III-A describes the design of the symbolic controller that is deployed as the upper-level controller; Section III-B elaborates on the velocity feedback controllers applied in the low-level velocity control loop.
@@ III-A Symbolic Control
Symbolic control [27] is an approach to derive provably safe controllers that satisfy complex logical specifications, including objectives and constraints in linear temporal logic. Symbolic control relies on the existence of a symbolic system, which is a finite abstraction of the continuous system. We use the approach in [18] to obtain a feedback refinement relation for the vessel, as modeled by (1), that enables abstraction-based controller synthesis and leads to the control loop seen in Fig. 2.
To better understand the correctness guarantees we achieve through this process, let $S_{1}$ and $S_{2}$ represent the plant (1) and an abstraction of $S_{1}$, respectively, such that they are related through a feedback refinement relation, $Q$. Further, let $\Sigma_{1}$ be a specification on the state and input space and assume that $\Sigma_{2}$ is the corresponding abstract specification. Then, $S_{2}$ contains the dynamical behaviors of $S_{1}$, and it is a finite-state model that is amenable to synthesis using standard search algorithms.
Given this $Q$, the following holds: if the abstract controller $C$ solves the abstract control problem $(S_{2},\Sigma_{2})$, then the refined controller $C\circ Q$ solves the actual control problem $(S_{1},\Sigma_{1})$ [Theorem III.5, [18]]. Therefore, the resulting symbolic controller, $C\circ Q$, is guaranteed to output a list of all possible control actions that satisfy the given specification. We optimize over this list to determine the optimal control input, resulting in a velocity command executed by a low-level velocity controller, as shown in Fig. 1.
@@
Figure 2: Closed loop: abstraction, refinement, and optimization.
We adopt the convention where the states of system $i$ are denoted with subscript $i$ and $[a,b]$ denotes a discrete interval. We aim to guide the vessel (1) throughout a region with boundary, $B$, into a target docking position, $T_{1}$, while avoiding obstacles $O^{1}_{1}$ and $O^{2}_{1}$ as seen in Fig. 3. We dynamically measure the state of the vessel and obstacles. Then we use this information to perform the controller synthesis task that results in new control inputs that are sent to the vessel every $2$ seconds. This update frequency accounts for the settling time of the low-level velocity controller. We formulate our high-level control objective $(S_{1},\Sigma_{1})$ where $S_{1}$ is a mathematical representation of the physical vessel model and $\Sigma_{1}$ is the following specification:
@@
with $U_{1}$ as the set of inputs and $X_{1}$ as the set of vessel states.
$O^{1}_{1}$$O^{1}_{2}$$O^{2}_{1}$$O^{2}_{2}$$T_{2}$$T_{1}$$B$
@@
$S_{2}$ is connected to the plant via a static quantizer, irrespective of the specification imposed on the plant. In defining $S_{2}$, we introduce margins for over-approximations of obstacles and an under-approximation for the target docking position. This can be seen in Fig. 3 as $O^{1}_{2}$, $O^{2}_{2}$, and $T_{2}$. These margins account for the geometrical shape of the vessel. Alternatively, they can be set according to other factors, e.g.  [28, 29]. Therefore, the abstract specification, $\Sigma_{2}$, of our system is (3) where we substitute $O^{1}_{1},O^{2}_{1},T_{1}$ with $O^{1}_{2},O^{2}_{2},T_{2}$, respectively.
@@
The abstract system, $S_{2}$, has an associated transition function among its states, as illustrated in Fig. 4. A standard search algorithm [27] is performed to determine safe control commands that allow the vessel to achieve its goal position. A list of all possible control actions that solve this problem are returned each time the synthesis task is performed with updated vessel and obstacle positions.
Once a list of safe possible actions is computed, the plant may be given any one of the control commands. Every action on this list will result in safe progress towards the plant’s goal. We optimize this selection process for realistic vessel behavior, using a cost function that favors forward motion and reduces the norm between sequential control actions to reduce fluctuations in the control reference signal. Let $\mathcal{A}=\{\sigma_{1},\sigma_{2},\dots,\sigma_{n}\}$ be a list of actions where $\sigma_{i}\in\mathbb{R}^{3}$, $\sigma_{pre}$ be the previously chosen action, and $\Delta\sigma_{i}:=\sigma_{i}-\sigma_{pre}$. We define the cost function $J(\sigma_{i})$ as
@@
where $W\in\mathbb{R}^{6\times 6}$ is positive semi-definite, as designed in Section IV-E; this matrix is application specific. At every time step, we select the control action that minimizes the cost, $\nu_{ref}:=\sigma^{*}=\text{arg}\min_{\sigma_{i}\in\mathcal{A}}J(\sigma_{i})$.
@@ III-B Velocity Feedback Controller
The velocity feedback controller, illustrated in Fig. 5, controls the vessel’s surge, sway, and yaw velocities, represented by $\nu_{ref}$, using a Multiple-Input Multiple-Output (MIMO) PID velocity controller, an extended Kalman filter (EKF) for state estimation, and a maneuvering based dynamic thrust allocation algorithm [30]. The PID controller calculates the desired force, $\tau_{cmd}$, to achieve the reference velocity $\nu_{ref}$. Then, the thrust allocation algorithm determines the necessary thruster forces, $F_{1},F_{2},F_{3}$, and azimuth angles, $\alpha_{1},\alpha_{2}$, to apply these commands. By performing reference tracking, the PID controller smooths the discretized reference signal sent by the symbolic controller. Position and velocity predictions are computed using the robot_localization package’s EKF implementation, as described in [31]. The measurement sources vary based on setup, as detailed in Section IV.